I wasn't sure what category to ask this in, but this one seemed like the best bet. My question involves force and surface area.

Let's say there are two objects, Object A and Object B:
-Object A weighs a little less than Object B
-Object A has far more surface area than Object B
-Both Object A and Object B are being supported by identical structures

Which will have more force on the structure? How do I calculate the force (or would it be pressure?) that each object has on the structure?
1
Expert's answer
2011-10-07T08:36:06-0400
Force acting onthe structure is pressure:
P=F/S

where F is theweight of an object, S - the surface of an object that touches the structure.
So
Fa<Fb ; Sa>>SbPa=Fa/Sa; Pb=Fb/Sb

So A has lessforce on the struct. as its surface is bigger and force smaller.